# Atomic Smart Contract Interoperability with High Efficiency via Cross-Chain Integrated Execution

### v1

@@ IV-C On-Chain Verification
Although off-chain clone and deployment can achieve efficient cross-chain logic deployment, it does not guarantee security as there may be malicious relayers present in the system. A malicious relayer could potentially modify the contract bytecode or deploy a wrong contract. To ensure verifiability and security during the off-chain clone and deployment process, we propose the on-chain cross-chain verification scheme to compare the cloned logic contract with the original contract and verify its correctness. The on-chain verification process is shown in Figure  3(b).
After the cross-chain deployment is completed, the cross-chain dApp provider can initiate cross-chain verification on the execution chain by calling the Verification() function of the bridging contract. The bridging contract will search the cloned contract bytecode of address $Addr^{\prime}_{\text{L}}$, and calculate the hash of the bytecode. Then the bridging contract triggers an event includes the hash computation result. The relayers are responsible for transmitting this information (a cross-chain transaction) along with its Merkle proof [52] to the invoked chain. In the invoked chain, the Merkle proof with the cross-chain transaction is first verified to ensure the result has already reached consensus on the execution chain. The bridging contract then searches the corresponding local contract bytecode based on the address $Addr_{\text{L}}$, and calculates the hash of the bytecode. Then, the bridging contract will verify whether it matches the hash transmitted across the chain. The result of the verification is then returned.
If the verification is successful, the cloned contract will be marked as verified and allowed for subsequent cross-chain invocations, and the relayer that performed the cross-chain deployment will be rewarded. If the verification fails, the relayer will be penalized, and the off-chain clone and deployment process will be restart.
@@ V-A Atomic Integrated Execution
Updating. The Updating process unlocks and updates all the invoked contract states on the invoked chains. After Integrated Execution is completed on the execution chain, the cross-chain dApp contract triggers an event to update the result via bridging contract. The relayers will distribute the result to the invoked chains. The bridging contract on each invoked chain verifies the Merkle proof of the cross-chain transaction and then invokes the UpdateState() function of each invoked contract, which will unlock and update the states of the invoked contracts.
@@
Figure 4: The CCSCI process in sequential invocation (left) and in IntegrateX’s Cross-Chain Smart Contract Integrated Execution Protocol (right).
Timeout. Additionally, to prevent the invoked contract state from being locked for an extended period, a design of timeout will be determined by the dApp developer within the application and managed by the execution chain. For transactions that time out, such as when the locking is not completed or execution is not finished for an extended period, the execution chain will mark the cross-chain call as failed and send a cross-chain event to unlock the relevant contracts.
Finality. In IntegrateX, all cross-chain transactions must wait until the consensus on the initiating chain is finalized (or, for Nakamoto-type consensus, highly likely to be finalized) before being committed to the target chain, to ensure cross-chain security.
@@ V-B Transaction Aggregation
A large amount of cross-chain transactions for transmitting state incurs significant gas consumption. As shown in Figure 4 left, handling CCSCI by sequential invocation may need to invoke contracts on the same chain in multiple rounds, which cause multiple rounds of cross-chain state transfers. To address this issue, we design a transaction aggregation mechanism to reduce gas costs caused by multiple invocations of different contracts and state transfers on the same chain.
As shown in Figure 4 right, our protocol locks all required states simultaneously, allowing multiple states on the same chain to be locked together, even when the calls are non-contiguous. The transaction aggregation mechanism combines all state requests on the same chain into a single transaction, reducing the number of cross-chain transactions. Similarly, during state updates, the transaction aggregation mechanism reduces the number of update requests. Therefore, for cross-chain calls involving multiple contracts on the same chain, this approach ensures that the number of cross-chain transactions is equal to the number of invoked chains (rather than the number of contracts), thereby reducing gas consumption.
@@ VII-A Implementation
We implement a prototype of IntegrateX based on a cross-chain communication project Bitxhub [53] (its transport layer is similar to IBC). The relayers are implemented in Golang, while the bridging contracts are implemented using Solidity. In addition, we implement several EVM-compatible blockchains in Golang leveraging go-ethereum clients. For performance comparison, we also implement a baseline CCSCI protocol GPACT [23]. As mentioned in the paper, GPACT is one of the most advanced existing protocols that can guarantee overall atomicity for complex CCSCI. However, GPACT has limitations in efficiency, as it requires multiple rounds of cross-chain execution and message exchange. Furthermore, we use Solidity to implement the Train-and-Hotel example mentioned in the paper, along with smart contracts in various other scenarios, to demonstrate the IntegrateX’s performance in real use cases and under different conditions.
@@
We test the performance of IntegrateX on multiple servers. Each server is equipped with an Intel® Core(TM) i5-10400F CPU @ 2.90 GHz and 15.5 GB RAM running 64-bit Ubuntu 22.04 LTS. We deploy 4 relayers via Docker containers on multiple servers. We also deploy 3 blockchains. Each blockchain contains 4 nodes and uses the Proof-of-Authority consensus algorithm. The block time is set to be 5 seconds. Additionally, the gas calculation method in the experiments is the same as that used in Ethereum. For every experiment, we select Blockchain 1 (bc1) as the execution chain in the IntegrateX system, with the other chains as invoked chains.
@@
VII-B Experimental Results
VII-B1 Performance under Train-and-Hotel Use Case
@@
As shown in Figure 5(a), IntegrateX decreases total latency by 51.2% compared to GPACT. On bc1 (execution chain), the latency of IntegrateX is reduced by 46.2% compared to GPACT. The reason is that, due to the integrated execution, IntegrateX could integrate and execute all the related execution logic within one block, avoiding multiple rounds of cross-chain execution during CCSCI. On bc2, IntegrateX also reduces the latency by 60% compared to GPACT. This is because the train contract is invoked multiple times, requiring multiple rounds of state transfers in GPACT, whereas in IntegrateX, only a single round of state transfer is needed.
@@
VII-B2 Performance under Different Invocation Complexity
We evaluate the performance of latency and gas consumption of IntegrateX and GPACT with different CCSCI complexity. We vary the call tree depth to represent different levels of complexity in cross-chain invocations.
@@ VII-B2 Performance under Different Invocation Complexity
(b) Gas consumption
Figure 6: Latency and gas consumption under different call tree depths.
@@
The gas consumption results are shown in Figure 6(b). Due to the lower invocation complexity, the gas savings from transaction aggregation are not significant, when the call tree depth is 3, the gas consumption for CCSCI in IntegrateX is nearly the same compared to GPACT. As the invocation complexity increases, when the call tree depth reaches 4, transaction aggregation reduces gas consumption in IntegrateX by 9.8% compared to GPACT. And with further increases in invocation complexity, transaction aggregation will reduces more gas consumption in IntegrateX. We also evaluate the gas consumption between IntegrateX and IntegrateX without transaction aggregation. The experimental results show that as the complexity of CCSCI increases, the transaction aggregation mechanism significantly reduces gas consumption in CCSCI. When the call tree depth is 3, the transaction aggregation mechanism reduces gas consumption by 5.4%, and when the call tree depth is 4, it reduces gas consumption by 19.3%. This suggests that the transaction aggregation mechanism achieves greater improvements in gas consumption as the complexity of CCSCI increases.
VII-B3 Concurrency Performance with Fine-Grained State Lock
We now measure the concurrency performance of IntegrateX with and without the fine-grained state lock mechanism. To show the concurrency performance, we initiate multiple cross-chain calls on the same state simultaneously and compare the time consumed by the cross-chain dApp. The experimental results are shown in the Figure 7.

### v2

@@ IV-C On-Chain Verification
Although off-chain clone and deployment can achieve efficient cross-chain logic deployment, it does not guarantee security as there may be malicious relayers present in the system. A malicious relayer could potentially modify the contract bytecode or deploy a wrong contract. To ensure verifiability and security during the off-chain clone and deployment process, we propose the on-chain cross-chain verification scheme to compare the cloned logic contract with the original contract and verify its correctness. The on-chain verification process is shown in Figure  3.
After the cross-chain deployment is completed, the cross-chain dApp provider can initiate cross-chain verification on the execution chain by calling the Verification() function of the bridging contract. The bridging contract will search the cloned contract bytecode of address $Addr^{\prime}_{\text{L}}$, and calculate the hash of the bytecode. Then the bridging contract triggers an event includes the hash computation result. The relayers are responsible for transmitting this information (a cross-chain transaction) along with its receipt proof and block header to the invoked chain. The receipt proof serves as a cryptographic proof that a specific event (e.g., a function execution or asset transfer) has indeed occurred on the source chain. It typically includes a Merkle proof [50], which attests that the corresponding event log is embedded in a finalized block. The block header provides the receiptsRoot, which serves as the cryptographic commitment to the receipt set of the block and is used to verify the inclusion of the event log via the receipt proof. This enables the invoked chain to verify the authenticity of the event without relying on trust assumptions about the relayer, thereby ensuring the integrity of cross-chain execution. In the invoked chain, the receipt proof with the cross-chain transaction is first verified to ensure the result has already reached consensus on the execution chain. The bridging contract then searches the corresponding local contract bytecode based on the address $Addr_{\text{L}}$, and calculates the hash of the bytecode. Then, the bridging contract will verify whether it matches the hash transmitted across the chain. The result of the verification is then returned.
The bridging contract on the execution chain receives the verification result from the invoked chain. To ensure authenticity, the execution chain first verifies that the provided block header originates from the invoked chain and is finalized (e.g., via the bridging contract). Then, using the accompanying receipt proof, it reconstructs the Merkle path to the target receipt and verifies that the computed receiptsRoot matches the one in the provided block header. This process guarantees the integrity and authenticity of the event data emitted on the invoked chain. If the result confirms that the bytecode of the cloned contract on the execution chain matches the original, the bridging contract marks the cloned contract as verified. Only verified contracts are permitted to participate in subsequent cross-chain invocations. Upon successful verification, the relayer responsible for the deployment is rewarded. If the verification fails, the relayer will be penalized, and the off-chain clone and deployment process will be restart.
@@ V Cross-Chain Smart Contract Integrated Execution Protocol
Figure 4: IntegrateX’s Cross-Chain Smart Contract Integrated Execution Protocol.
@@ V-A Atomic Integrated Execution
Updating. The Updating process unlocks and updates all the invoked contract states on the invoked chains. After Integrated Execution is completed on the execution chain, the cross-chain dApp contract triggers an event to update the result via bridging contract. The relayers will distribute the result to the invoked chains. The bridging contract on each invoked chain verifies the Merkle proof of the cross-chain transaction and then invokes the UpdateState() function of each invoked contract, which will unlock and update the states of the invoked contracts. It is important to note that in IntegrateX, the address of the bridging contract is a required parameter during the initialization of a state contract. In the state contract, an address variable is used to store the address of the bridging contract. Before executing any function related to state locking or updating, the contract verifies that the caller is the registered bridging contract. Only upon successful verification are the locking or updating operations permitted. This design allows the state contract to grant the bridging contract the permission to invoke specific internal functions, thereby enabling it to accept state locking and state updates initiated by the bridging contract.
@@
For security concerns, within CCSCI, only the bridging contract is authorized (configured in the contract logic) to unlock the state contracts, while the relayer has no direct access to unlock any locked contract state. Furthermore, the relayer is incapable of forging an unlock request to the bridging contract, since any maliciously crafted message cannot pass the Merkle proof verification and will therefore be rejected during contract execution. This design ensures the security and integrity of CCSCI execution. Besides, the rollback process does not involve any state updates. All pending state changes reside on the execution chain, and rollback merely discards the result of the integrated execution on the execution chain while releasing the locked states on the invoked chains. As such, IntegrateX does not risk unintended consequences for other CCSCI.
Timeout. Additionally, to prevent the invoked contract state from being locked for an extended period, a design of timeout will be determined by the dApp developer within the application and managed by the execution chain. Timeout is a widely adopted mechanism in lock-based solutions [25, 22]. In IntegrateX, both the bridging contract (set by our protocol) and the developer-defined smart contract can specify a maximum timeout. During execution, the effective timeout is determined by taking the minimum of these two values, which prevents the dApp provider from setting an indefinite timeout period. Moreover, timeout in blockchain systems is typically measured in terms of block depth rather than wall-clock time. Therefore, although the actual block time may vary across different blockchains, the maximum timeout in IntegrateX can be aligned approximately by using chain-specific block depths. For example, blockchains such as Ethereum[2] tend to have longer block intervals (e.g.,  12 seconds), so a timeout of 10 blocks would translate to roughly 2 minutes. In contrast, blockchains such as BNB Smart Chain [51] may have shorter block times (e.g., 2-3 seconds), in which case a timeout of 10 blocks would correspond to a much shorter wall-clock duration. Moreover, the timeout is typically set to exceed the finality confirmation time of the underlying blockchain for safety. For transactions that time out, such as when the locking is not completed or execution is not finished for an extended period, the execution chain will mark the cross-chain call as failed and send a cross-chain event to unlock the relevant contracts.
Finality. In IntegrateX, all cross-chain transactions must wait until the consensus on the initiating chain is finalized (or, for Nakamoto-type consensus [1], highly likely to be finalized) before being committed to the execution chain or invoked chain, to ensure cross-chain security. For different blockchains, IntegrateX allows the configuration of varying confirmation depths to ensure that the result of a transaction is sufficiently finalized. Specifically, it supports setting a block confirmation threshold in the bridging contract, enabling tailored finality strategies for heterogeneous blockchains.
To ensure that only finalized states are used on the execution/invoked chain, the bridging contract verifies the confirmation status of the submitted block header from the invoked/execution chain. It compares the submitted header with the latest block height of the invoked/execution chain and ensures that the block has reached the required confirmation depth. If this condition is not satisfied, the cross-chain message will be rejected (can be resend). In addition, the bridging contract checks the validity of the message through the receipt proof. Only when both the confirmation threshold and receipt proof verification succeed will the event be accepted and its associated state be used for further execution. For different blockchains, the threshold setting is typically related to how long it takes for the blockchain itself to reach finality. Information regarding finality is usually defined in their project documentations. IntegrateX only needs to set the corresponding threshold in the bridging contract according to the definition in their documentations before connecting to their blockchains. This flexible confirmation mechanism ensures that cross-chain execution is performed based on finalized and reliable states.
@@ V-B Transaction Aggregation
(a) Invocation without TA
(b) Invocation with TA
Figure 5: An example of IntegrateX’s Transaction Aggregation mechanism.
A large amount of cross-chain transactions for transmitting state incurs significant gas consumption. Handling CCSCI by sequential invocation may need to invoke contracts on the same chain in multiple rounds, which cause multiple rounds of cross-chain state transfers. To address this issue, we design a transaction aggregation mechanism to reduce gas costs caused by multiple invocations of different contracts and state transfers on the same chain.
Our protocol locks all required states simultaneously, allowing multiple states on the same chain to be locked together, even when the calls are non-contiguous. The transaction aggregation mechanism combines all state requests on the same chain into a single transaction, reducing the number of cross-chain transactions. Similarly, during state updates, the transaction aggregation mechanism reduces the number of update requests. Therefore, for cross-chain calls involving multiple contracts on the same chain, this approach ensures that the number of cross-chain transactions is equal to the number of invoked chains (rather than the number of contracts), thereby reducing gas consumption.
Figure 5 shows an example that the execution chain needs to lock the state of both Contract A and Contract B on the invoked chain. Sequential invocation needs to invoke Contract A and Contract B in multiple rounds, as illustrated in Figure 5(a). By using transaction aggregation mechanism, the execution chain only needs to send a single transaction to lock the state of both Contract A and Contract B, as shown in Figure 5(b). Through this mechanism, the execution chain can avoid multiple rounds of cross-chain messaging, improving efficiency. Furthermore, the reduction in the number of transactions also leads to lower gas consumption.
@@ V-C Fine-Grained State Lock
For example, as shown in Listing 1 in Section IV-A1, users can book hotel rooms through a smart contract. For simplicity, we assume all rooms are of the same type, and the remain variable represents the number of available rooms. Without fine-grained state locking, a booking attempt by one user would lock the entire contract state, preventing concurrent transactions. In the SHotel contract, a fine-grained locking mechanism is implemented using a lock_bag structure, which allows partial locking of the remain variable. This enables other transactions to access and interact with the remaining state. During the locking period, a lockpool temporarily holds the locked portion of the state. Developers can configure the lock_size parameter to determine how much of the state is locked per transaction, helping to prevent failures caused by insufficient locked state.
@@ VI-A Security in Hybrid Cross-Chain Smart Contract Deployment Protocol
It is worth noting that the concept of a relayer is not originally proposed by IntegrateX, nor is the assumption of at least one honest relayer unique to this system. In fact, all existing cross-chain protocols rely on the presence of at least one genuinely honest and non-compromised relayer [21, 22, 23, 36, 39]. This relayer must not be forgeable or impersonated by an adversary; otherwise, the threat model would be fundamentally violated and the protocol’s security guarantees would no longer hold. During cross-chain communication, the relayer solely acts as a message carrier. It listens to on-chain events and forwards the corresponding message along with a receipt proof (mentioned in Section IV-C), which attests to the inclusion of the event in the block header of the source chain. Upon receiving the message, the invoked chain verifies the receipt proof to ensure the message was indeed generated by the execution chain. Only after successful verification does the bridging contract on the invoked chain reconstruct and execute the intended contract call. This verification mechanism is symmetric and also applies to messages sent from the invoked chain back to the execution chain, thereby ensuring the integrity of cross-chain communication regardless of relayer trustworthiness. Therefore, when inconsistencies arise due to a relayer maliciously modifying the transmitted message, the tampered data will be rejected by the invoked chain as it fails the on-chain proof verification. Only messages that pass the verification will be accepted.
Furthermore, in the case of blockchain forks leading to conflicting relayer messages, IntegrateX defers execution until one fork becomes finalized (mentioned in Section V-A). Once finality is reached, the corresponding transaction results are adopted for further processing.
@@ VI-B Security in Cross-Chain Smart Contract Integrated Execution Protocol
It is important to note that, similar to existing cross-chain solutions, IntegrateX incorporates an acknowledgment (receipt) mechanism for cross-chain messaging [52, 22]. In the Updating phase of integrated execution, the execution chain receives receipts from all invoked chains. The final result is output on the execution chain only after all receipts confirm that the state updates have been successfully completed on the respective chains. This mechanism ensures that the execution chain is informed whether the updating states transaction has been successfully executed. Only upon successful execution will the system proceed to the next step. Any failure in the cross-chain execution triggers a full rollback of the transaction, thereby preserving the overall atomicity of IntegrateX. As a result, even if an update operation on the invoked chain fails (which barely happens), the entire transaction will be rolled back, maintaining the system’s atomicity guarantees.
@@ VII-A Implementation
We implement a prototype of IntegrateX based on a cross-chain communication project Bitxhub [52] (its transport layer is similar to IBC). The relayers are implemented in Golang, while the bridging contracts are implemented using Solidity. For the underlying blockchain, we do not rely on a custom-built chain. Instead, we leverage the existing go-ethereum client to implement several EVM-compatible blockchains in Golang, without making any modifications to the underlying blockchain infrastructure. For performance comparison, we also implement a baseline CCSCI protocol GPACT [21]. As mentioned in the paper, GPACT is one of the most advanced existing protocols that can guarantee overall atomicity for complex CCSCI. However, GPACT has limitations in efficiency, as it requires multiple rounds of cross-chain execution and message exchange. Furthermore, we use Solidity to implement the Train-and-Hotel example mentioned in the paper, along with smart contracts in various other scenarios, to demonstrate the IntegrateX’s performance in real use cases and under different conditions.
For the ease of implementation and experimentation, the underlying blockchain employs a Proof-of-Authority (PoA) mechanism for leader selection, and achieves consensus through Practical Byzantine Fault Tolerance (PBFT). PoA is a consensus protocol in which a limited set of pre-approved validators, referred to as authorities, are responsible for block production.
To support our proposed protocols, we implement a bridging contract on the blockchain. The protocol requires no modification to the underlying blockchain infrastructure; it only involves deploying these bridging contracts. We implement the functions regServer and regState to register the logic contract and the state contract, respectively. During the on-chain verification process, the compareBytes function in the bridging contract is used to compare the bytecode of the cloned contract with that of the original contract on the invoked chain, ensuring that the contract has been correctly cloned. In the Cross-Chain Smart Contract Integrated Execution Protocol, the bridging contract is also responsible for issuing the CCSCI transaction. To this end, we implement the lockState and updateState functions to lock and update the contract state on the invoked chain. Importantly, this integration is lightweight and non-intrusive, requiring no changes to the underlying blockchain beyond smart contract deployment.
Furthermore, IntegrateX supports cross-chain interoperability via a relayer built upon BitXHub’s cross-chain communication framework. The relayer includes a receipt mechanism to ensure the reliability and traceability of message delivery. Unlike traditional cross-chain systems that rely on trusted relayers for message validation, the relayer in IntegrateX is entirely trustless. Its sole responsibility is to relay raw cross-chain data—such as events, receipts, and Merkle proofs—between chains. All critical verification, including block header validation, Merkle proof checking, and event authenticity verification, is executed on-chain through smart contracts. This design eliminates the need to trust any off-chain component and ensures that incorrect or forged messages submitted by a malicious relayer will be rejected by the verification logic on the execution chain. Consequently, the relayer can be treated as a permissionless data courier, significantly enhancing the security and decentralization of the system.
VII-B Experimental Setup
@@
We test the performance of IntegrateX on multiple servers. Each server is equipped with an Intel® Core(TM) i5-10400F CPU @ 2.90 GHz and 15.5 GB RAM running 64-bit Ubuntu 22.04 LTS. We deploy 4 relayers via Docker containers on multiple servers. We also deploy 3 blockchains, each with 4 nodes. To decide block time, we notice that existing well-known EVM-compatible blockchains (e.g., BNB Smart Chain [51], Polygon [53], Optimism [54], Ethereum [2]) typically adopt block times ranging from 2 to 12 seconds. As a result, the default block time in the experiments is set to be 5 seconds to maintain generality while ensuring a reasonable balance between performance and realism. We set each block to contain up to 4096 transactions. Additionally, the gas calculation method in the experiments is the same as that used in Ethereum. For every experiment, we select Blockchain 1 (bc1) as the execution chain in the IntegrateX system, with the other chains as invoked chains.
@@
VII-C Experimental Results
VII-C1 Performance under Train-and-Hotel Use Case
@@
IntegrateX decreases total latency by 51.2% compared to GPACT, as shown in Figure 6(a),. On bc1 (execution chain), the latency of IntegrateX is reduced by 46.2% compared to GPACT. The reason is that, due to the integrated execution, IntegrateX could integrate and execute all the related execution logic within one block, avoiding multiple rounds of cross-chain execution during CCSCI. On bc2, IntegrateX also reduces the latency by 60% compared to GPACT. This is because the train contract is invoked multiple times, requiring multiple rounds of state transfers in GPACT, whereas in IntegrateX, only a single round of state transfer is needed.
@@
Figure 7: Latency of handling CCSCI in IntegrateX with different block times.
Current EVM-compatible blockchains exhibit a range of block times. The results presented in Figure 7 show the experiments conducted on the Train-and-Hotel use case under different block time settings. The results indicate that IntegrateX consistently demonstrates significantly lower latency compared to GPACT across all block times. This improvement is primarily due to IntegrateX’s ability to avoid multiple rounds of cross-chain execution during CCSCI, thereby reducing overall latency. As the block time increases, the latency of both IntegrateX and GPACT exhibits an approximately linear growth trend.
VII-C2 Performance under Different Invocation Complexity
We evaluate the performance of latency, throughput and gas consumption of IntegrateX and GPACT with different CCSCI complexity. We vary the call tree depth to represent different levels of complexity in cross-chain invocations.
@@ VII-C2 Performance under Different Invocation Complexity
(b) Throughput
(c) Gas consumption
Figure 8: Latency, gas consumption and throughput under different call tree depths.
@@
The throughput results are shown in Figure 8(b). IntegrateX achieves significantly higher throughput than GPACT across varying call tree depths. Specifically, when the call tree depth is 2, 3, and 4, IntegrateX outperforms GPACT by 30.4%, 104.3%, and 199.1%, respectively. This improvement stems from IntegrateX’s transaction aggregation mechanism, which ensures that the number of transactions required to execute CCSCI remains constant, regardless of call tree depth. As a result, the throughput of IntegrateX remains nearly stable. In contrast, IntegrateX without transaction aggregation, as well as GPACT, both exhibit a linear increase in the number of required transactions as the call tree depth increases, leading to a corresponding linear decrease in throughput. Moreover, since integrated execution in IntegrateX requires fewer transactions than GPACT even without aggregation, it consistently achieves better throughput. Therefore, the performance advantage of IntegrateX becomes increasingly pronounced with deeper call trees.
The gas consumption results are shown in Figure 8(c). Due to the lower invocation complexity, the gas savings from transaction aggregation are not significant, when the call tree depth is 3, the gas consumption for CCSCI in IntegrateX is nearly the same compared to GPACT. As the invocation complexity increases, when the call tree depth reaches 4, transaction aggregation reduces gas consumption in IntegrateX by 9.8% compared to GPACT. And with further increases in invocation complexity, transaction aggregation will reduces more gas consumption in IntegrateX. We also evaluate the gas consumption between IntegrateX and IntegrateX without transaction aggregation. The experimental results show that as the complexity of CCSCI increases, the transaction aggregation mechanism significantly reduces gas consumption in CCSCI. When the call tree depth is 3, the transaction aggregation mechanism reduces gas consumption by 5.4%, and when the call tree depth is 4, it reduces gas consumption by 19.3%. This suggests that the transaction aggregation mechanism achieves greater improvements in gas consumption as the complexity of CCSCI increases.
VII-C3 Concurrency Performance with Fine-Grained State Lock
We now measure the concurrency performance of IntegrateX with and without the fine-grained state lock mechanism. Concurrency refers to the system’s ability to execute multiple full cross-chain smart contract invocations in parallel, rather than isolated operations on a single chain. To show the concurrency performance, we initiate multiple cross-chain calls on the same state simultaneously and compare the time consumed by the cross-chain dApp. The experimental results are shown in the Figure 9.